OpenStudy (anonymous):

Simplify. y ^5 ÷ y ^3 · y ^2 y=________

OpenStudy (bahrom7893):

=y

OpenStudy (anonymous):

sorry made a mistake

OpenStudy (bahrom7893):

so what's the problem?

OpenStudy (anonymous):

i think it nvm

OpenStudy (anonymous):

i think there is some probs with the question.accordin to the question the answer should be 0 i think

OpenStudy (anonymous):

Simplify. y ^5 ÷ y ^3 · y ^2 y^___

OpenStudy (anonymous):

y^4

OpenStudy (bahrom7893):

y^(5-3+2) = y^0 = 1

OpenStudy (bahrom7893):

woops sorry dipin was right it is y^4

OpenStudy (bahrom7893):

lol i've always had trouble with arithmetic..

OpenStudy (anonymous):

ok i read your inf o it seid that you wonted to go to law school me to i can talk my way like a lawer all the time so i am going to law school

OpenStudy (anonymous):

but any was back to math